### v3

# High-Quality Mask Tuning Matters for Open-Vocabulary Segmentation

Abstract
Open-vocabulary image segmentation has been advanced through the synergy between mask generators and vision-language models like Contrastive Language-Image Pre-training (CLIP). Previous approaches focus on generating masks while aligning mask features with text embeddings during training. In this paper, we observe that relying on generated low-quality masks can weaken the alignment of vision and language in regional representations. This motivates us to present a new fine-tuning framework, named MaskCLIP++, which uses ground-truth masks instead of generated masks to enhance the mask classification capability of CLIP. Due to the limited diversity of image segmentation datasets with mask annotations, we propose incorporating a consistency alignment principle during fine-tuning, which alleviates categorical bias toward the fine-tuning dataset. After low-cost fine-tuning, MaskCLIP++ significantly improves the mask classification performance on multi-domain datasets. Combining with the mask generator in previous state-of-the-art mask-based open vocabulary segmentation methods, we achieve performance improvements of +1.7, +2.3, +2.1, +3.1, and +0.3 mIoU on the A-847, PC-459, A-150, PC-59, and PAS-20 datasets, respectively. Code is avaliable at https://github.com/HVision-NKU/MaskCLIPpp.
1 Introduction
(a) Performance: generated masks v.s. ground truth masks.
(b) Performance: CLIP classification v.s. ground truth classification.
Figure 1: Observations: 1(a) demonstrates the potential negative impact of low-quality generated masks on CLIP’s mask classification learning, while 1(b) showcases the untapped generalization potential of existing mask generators. Mask generators are trained on the COCO [36], with results reported on the ADE20K [80].
Image segmentation is one of the most extensively studied tasks in computer vision, which aims to partition an image into several regions where each region corresponds to the same object or shares consistent semantics. Traditional image segmentation models are often defined on a closed vocabulary. When new classes need to be segmented, it typically requires redefining the vocabulary set, annotating data, and retraining the model. However, the cost of annotating image segmentation on a large vocabulary is high. Large-scale image-text pre-training models, such as CLIP [47] and ALIGN [25], have demonstrated strong zero-shot recognition capabilities, sparking interest in transferring these abilities to image segmentation tasks. Several studies [15, 26, 76, 27] have explored how to adapt mask generators [9, 10] and pre-trained CLIP to each other to achieve open-vocabulary segmentation at both the semantic and instance levels. However, our analysis based on two key observations reveals critical shortcomings in existing research that have been consistently overlooked.
First, we establish a baseline method enabling CLIP to perceive mask categories by referencing prior methodologies [73, 76]. To evaluate the impact of mask quality on inference performance, we conduct comparative experiments using generated versus ground-truth masks across various pre-trained CLIP models. As shown in Fig. 1(a), generated masks exhibit inferior performance compared to ground-truth masks due to their low segmentation accuracy, semantic ambiguity, and excessive redundancy. Our experiments further demonstrate that employing higher-quality masks during training consistently yields superior performance under identical inference protocols.
Second, to validate the strong generalization capability of existing mask generators, we evaluate multiple architectures pre-trained on the COCO dataset [36] and evaluate on the ADE20K dataset [80] with a different vocabulary set. Through IoU-based bipartite matching, we reassign “ground-truth categories” to predicted masks to determine whether these generators produce masks for unseen categories. As illustrated in Fig. 1(b), the performance could be substantially improved if mask classification were perfect. This discrepancy suggests that existing mask generators inherently produce numerous valid masks for unseen categories, but their potential is severely limited by erroneous mask classification.
Based on these observations, we propose to eliminate the reliance on mask generators during training and directly fine-tune CLIP using high-quality image segmentation data. However, since image segmentation data is typically defined on closed vocabularies, mitigating the issue of CLIP overfitting the training data after fine-tuning poses an important challenge. Existing approaches either integrate a frozen CLIP during inference [72, 71, 76] or employ a frozen teacher model for distillation [26, 27]. The former may limit the model’s performance improvement, while the latter often requires careful tuning of distillation strategies and hyperparameters. To address this issue more effectively, we analyze the causes of modality misalignment in prior works and propose that the introduction of additional parameters should adhere to the consistent alignment principle, thereby preserving CLIP’s original embedding space throughout the alignment optimization process. The outcome of our explorations is a new CLIP fine-tuning framework, coined as MaskCLIP++.
Experimental results demonstrate that MaskCLIP++ can significantly enhance CLIP’s masked classification capability across datasets spanning diverse domains. Additionally, MaskCLIP++ can be used with different mask generators to achieve open-vocabulary segmentation at the semantic- or instance-level. Compared to recent state-of-the-art methods, MaskCLIP++ offers lower training cost in terms of training time and memory usage, while achieving even better performance. We hope that the universality and efficiency of MaskCLIP++ could inspire further research in open-vocabulary segmentation.
2 Related Work
Mask-based segmentation. Mask-based segmentation is a technique that views image segmentation as the process of mask generation and mask classification, which is also known as region-based segmentation. To address instance and panoptic segmentation, [22, 62, 60, 79] proposed various mask generators. Recently, [9] unified the semantic, panoptic and instance segmentation into a single architecture. MaskFormer [9] uses learnable queries to interact with image features, which are then decoded as masks and class scores. Mask2Former [10] further accelerates the convergence of MaskFormer and reduces memory usage through multi-scale decoder design, using masks from shallow decoders as priors for deeper layers, and employing sampling-based optimization. Current mask-based OVS methods need almost to retrain an open-vocabulary mask generator based on Mask2Former [74, 71, 76]. In contrast, our method can directly use the Mask2Former trained on a closed vocabulary during inference, reducing training cost and increasing flexibility in usage.
(a) Adapt Mask Generator to CLIP
(b) Adapt CLIP to Mask Generator
(c) MaskCLIP++
Figure 2: Comparison of training pipeline between previous mask-based OVS methods. $\mathcal{L}$ denotes the loss function. 2(a) adapts the mask generator to CLIP and avoids overfitting by freezing CLIP. 2(b) adapts CLIP to the mask generator and avoids overfitting by distillation. “T” and “S” denote the teacher and student models, respectively. Our method 2(c) abandons the mask generator and avoids overfitting by consistency alignment.
Open-vocabulary segmentation. Depending on the data used, existing methods have developed into different paradigms. The first paradigm offers many heuristic solutions based on observations of CLIP without additional training [81, 59, 33, 55, 29, 30]. The second paradigm utilizes image-text pair data [57, 7] to train OVS models under weak supervision [69, 40, 70, 78, 44, 6, 39]. Due to the lack of precise segmentation locations, these methods suffer from poor segmentation quality. The third paradigm enhances segmentation quality by incorporating image-mask data [28] and knowledge from SAM [63, 77]. However, the use of SAM not only increases inference cost but also relies on users or other detectors to provide high-quality prompts. The fourth paradigm seeks to achieve mutual adaptation between the segmentation model and CLIP through fully supervised training on image segmentation datasets containing image, mask, and category triplets, thereby producing an open-vocabulary segmentation model. Based on the type of the segmentation model, these methods can be categorized into mask-based approaches [14, 72, 19, 20, 35, 46, 26, 15, 74, 73, 71, 76, 38, 27] and pixel/patch-based ones [31, 82, 34, 11, 68]. Pixel-based OVS methods typically focus on semantic segmentation and cannot perform instance-level segmentation independently. Additionally, some unique approaches to achieving OVS involve image retrieval [52, 1], feature distillation [8, 65] or model merging [61]. Some unified large models also incorporate the functionality of OVS [83, 51]. This paper focuses on efficiently transferring CLIP on limited segmentation datasets to improve its performance on semantic- and instance-level open-vocabulary segmentation tasks.
3 Method
3.1 Revisiting Image-Text Embeddings Alignment for OVS
We first revisit the roles of the visual encoder (CLIP-V) and text encoder (CLIP-T) in the pretrained CLIP model. Denote the image embeddings from CLIP-V as $E_{i}\in\mathbb{R}^{1\times D}$, where $D$ represents the dimension of embedding space. To create text embeddings, a vocabulary set of $K$ categories is transformed into the corresponding sentences using templates like “A photo of {}”, and then encoded by CLIP-T, yielding $E_{t}\in\mathbb{R}^{K\times D}$. The original CLIP model maps both images and text into an aligned embedding space. In this space, image-text pairs maintain semantic alignment, which we describe as “pre-alignment” and can often be measured using cosine similarity $\langle E_{i},E_{t}\rangle$. In open-vocabulary segmentation, preserving the pre-alignment property is crucial for adapting CLIP to zero-shot mask classification, as it enhances the alignment of mask and text embeddings.
Previous mask-based OVS methods obtain pre-aligned mask embeddings and text embeddings in two ways. As illustrated in Fig. 2(a), the first way aims to train a mask generator that adapts to CLIP, with the generalization mainly derived from the original frozen parameters of CLIP [73, 15]. However, attempts to align the embedding spaces of the mask generator and CLIP often fail [72, 76]. As illustrated in Fig. 2(b), the second way make CLIP sensitive to the quality of generated masks during training to enhance performance and avoids overfitting through distillation [26, 27].
We identify two limitations of these two types of methods. First, they rely on noisy generated masks, which can degrade the quality of the region representations (as shown by the vertical comparisons in Fig. 1(b)). This degradation can weaken the alignment between region-specific image and text embeddings. Second, there is a lack of an intuitive explanation and a straightforward solution for learning better alignment without severe overfitting. Moreover, freezing CLIP limits its transferability from global to local features, while distillation typically requires carefully designed strategies and incurs additional training cost.
To overcome these limitations, we propose a new training pipeline for OVS, named MaskCLIP++, which, to our knowledge, is the first method that attempts to leverage ground truth (GT) masks to obtain region features instead of relying on the mask generator during training. Furthermore, we propose the principle named “consistency alignment” to alleviate overfitting when adding extra parameters to optimize alignment. We will give detailed descriptions in what follows.
Figure 3: Detailed framework of MaskCLIP++ for OVS tasks. The PSM represents the parameterized similarity modeling, which is designed under the principle of consistency alignment. The mask generator is only used during inference, and can be flexibly replaced.
3.2 A New Fine-tuning Framework for CLIP
In this subsection, we detail how to use our framework to fine-tune CLIP and integrate it with an off-the-shelf mask generator. The overall framework is illustrated in Fig. 3.
Extracting mask embeddings. Taking CLIP-V with an $L$-layer Transformers as an example, inspired by SAN [73], we posit that the layers from $1$ to $K$ predominantly perform feature extraction (denoted as the Extractor), while layers from $K+1$ to $L$ primarily conduct spatial fusion (referred to as the Fuser).
For the $N$ image tokens $F^{(l)}\in\mathbb{R}^{N\times C}$ that input to the $l$-th layer of the fusion module, we introduce $Q$ flattened image masks $M\in[0,1]^{Q\times N}$ as conditions to transform the fusion scope from global to local. As formally delineated in Eq. 1, we compute mask-conditioned modulation weights $\phi^{l}$ to perform weighted aggregation on $F^{(l)}$, followed by the layer-wise channel projection operator $\mathrm{Proj}^{(l)}$, thereby generating mask-specific tokens $E_{m}^{(l+1)}\in\mathbb{R}^{Q\times C}$.
$E_{m}^{(l+1)}=\operatorname{Proj}^{(l)}\left(\phi^{(l)}(M)\cdot F^{(l)}\right),~{}\text{where}~{}K<l<L.$ (1)
By leveraging the input projections $q$ and $k$ from the original attention module, we incorporate the mask into the weighting function through Eq. 2, thereby constraining the weight distribution within the mask’s active regions.
$\phi^{(l)}(M)=\operatorname{softmax}\left(\frac{q(E_{m}^{(l)})k(F^{(l)})^{T}}{\sqrt{d}}+\alpha M_{\text{th}}\right),$ (2)
where $\sqrt{d}$ is the attention scaling factor, $\alpha\in(0,+\infty)$ is an learnable parameter, $M_{\text{th}}$ is the result of setting all values in $M$ that are less than $\max(M)/2$ to $\infty$- ∞.
When initializing $E_{m}^{(K+1)}$ with the CLS token in layer $K+1$, $E_{m}^{(L)}$ and $E_{t}$ exhibit pre-aligned properties. The first two rows of Tab. 1 quantitatively demonstrate the mask classification performance of pre-trained CLIP under this initialization paradigm.
Fig. 4 visualizes the normalized weight function $\phi$, which reveals the contributions of image tokens at different spatial positions to the generation of the current mask representation. After fine-tuning, more image tokens are incorporated into the construction of their corresponding mask representations.
Figure 4: Visualization of the weight function. The regions of interest are enclosed by yellow contours, and the deeper the red color at a position, the higher its importance to the region.
Parameterized similarity modeling with consistency alignment. Given the limited capacity of the original CLIP model in characterizing localized regions, a straightforward approach involves fine-tuning the original CLIP parameters alongside additional learnable parameters to reformulate the similarity metric between $E_{m}$ and $E_{t}$. We introduce a class of methods termed Parameterized similarity modeling (PSM) for this purpose.
However, when incorporating the mask generator into similarity modeling, the mask embeddings produced by the mask generator align almost exclusively with CLIP’s text embeddings on the training set, resulting in the misalignment illustrated in Fig. 2(a). We first analyze the cause of this overfitting to develop our similarity modeling approach.
To facilitate discussion, we simplify this approach as follows. Let $t\in\mathbb{R}^{D}$ denote a normalized text embedding from CLIP, and let $m_{1},m_{2}\in\mathbb{R}^{D}$ denote different mask embeddings. The initial similarity is defined as $s=m^{T}t$. Now, before computing the inner product across modalities, the mask embeddings are updated using a parameter $\Theta\in\mathbb{R}^{D\times D}$, resulting in a new similarity $r=(\Theta m)^{T}t$. Although $r$ can be optimized to be close to the target, an unintuitive phenomenon may arise: After mask embeddings are updated via gradient descent, the order of $r_{1}$ and $r_{2}$ is closer to the target, but the order of $s_{1}$ and $s_{2}$ may deviate further from the target. This phenomenon is illustrated by a toy example in Sec. A.2 in supplementary material.
Consequently, the original similarity relationships between CLIP’s embeddings are disrupted. Our experiments demonstrate that the additional parameter $\Theta$ fails to achieve CLIP’s level of generalization when fine-tuning on limited image segmentation data, resulting in overfitting. Building upon the aforementioned analysis, we identify inconsistent alignment phenomena in PSM that prone to overfitting. This implies that preserving CLIP’s original similarity ordering to the greatest extent possible during optimization enables more effective utilization of additional parameters for alignment enhancement. We formalize this critical finding as the consistency alignment principle in PSM.
Specifically, for all mask embeddings $E_{m}\in\mathbb{R}^{Q\times C}$ extracted from an image and the text embeddings $E_{t}\in\mathbb{R}^{K\times C}$ of $K$ dataset categories, their similarity matrix $S=E_{m}E_{t}^{T}$ was projected through an over-parameterized linear projection along a newly created orthogonal dimension relative to both $Q$ and $K$ dimensions. Finally, cross-entropy loss was computed based on category annotations of each mask, guided by the optimized similarity map.
Integrating with off-the-shelf mask generators. As the training process is entirely decoupled from the mask generator, we can flexibly employ various types of mask generators during inference to achieve semantic- or instance-level open-vocabulary segmentation.
Compared to previous OVS methods, MaskCLIP++ offers the following several advantages. First, we use ground-truth masks to directly fine-tune CLIP during training, which are less noisy, thus improving alignment quality and reducing training cost. Second, from the perspective of avoiding overfitting, the consistency alignment principle provides more optimization space than freezing CLIP [73, 76], and does not rely on carefully designed distillation strategies [65, 27], resulting in faster convergence with a single optimization objective.
4 Experiments
4.1 Experimental Setup
Baseline. In this section, we employ pre-trained models from EVA02 CLIP [54] with ViT architecture. Following the framework outlined in Sec. 3.2, we configure $L-K=2$ with $\alpha$ initialized to $e^{-5}$. The second row of Tab. 1 demonstrates the mask classification accuracy of this baseline implementation.
For open-vocabulary image segmentation evaluation, we adopt FC-CLIP (corresponding to Fig. 2(a)) and MAFT+ (corresponding to Fig. 2(b)) as baseline approaches. Their baseline performance metrics are documented in Tab. 2 and Tab. 3. By retaining the mask generators from both frameworks while replacing the mask classification component with MaskCLIP++, we demonstrate the performance enhancement in overall OVS capabilities through improved mask classification accuracy.
Datasets and Metrics. Aligning with established OVS methodologies, we fine-tune our model on the COCO Stuff dataset [5] with its fixed 171-category vocabulary and evaluate on datasets with different vocabularies: ADE20K (150-class and 847-class taxonomies) [80], Pascal Context (459-class and 59-class taxonomies) [43], and PASCAL VOC (20 foreground categories excluding background) [17]. The mean Intersection over Union (mIoU) metric serves as our primary evaluation criterion for OVS performance. Furthermore, we quantitatively demonstrate performance enhancements in both panoptic and instance segmentation tasks on ADE20K through established metrics: panoptic quality (PQ), segmentation quality (SQ), recognition quality (RQ) and mean average precision (mAP).
To demonstrate the generalizable improvement of MaskCLIP++ in mask classification in various scenarios, we perform evaluations on datasets from the MESS benchmark [4], where the accuracy of the top-1 ground truth mask (maskAcc) serves as the primary metric to quantify the capability of zero shot mask classification.
Implementation Details. Our implementation is built on the Detectron2 framework [67]. During training, the number of image tokens $N=32^{2}$. The AdamW optimizer is employed with a batch size of 4 for 20K iterations. The initial learning rate is set to $10^{-4}$ and decayed using cosine annealing. Following the setting of CAT-Seg [11], we only train the linear projection parameters of the query and value in the CLIP attention modules, and set their learning rate $10^{2}$ times the overall learning rate. The over-parameterized dimension of the linear layer in PSM is set to 768.
During inference, while maintaining the original aspect ratio, images are resized with shorter edges scaled to 392 and 800 pixels respectively for CLIP-V and Mask Generator inputs. Similar to previous works [71, 76], we also ensemble the class scores of the mask generator and MaskCLIP++ on the training categories of the mask generator.
A-847 PC-459 A-150 PC-59 Stuff Citys General Earth Medical Engineer Agriculture
Pretrained
OpenAI CLIP [47] 32.0 44.8 51.0 57.9 46.2 46.5 54.2 61.1 61.0 36.7 51.0
EVA02 CLIP [54] 35.2 44.8 52.7 54.6 45.0 44.9 56.9 60.5 61.7 33.8 52.4
Finetuned on COCO
CLIPSelf [65] 33.6 49.0 56.1 67.5 50.6 52.1 52.5 45.7 55.5 45.3 56.3
CAT-Seg [11] 33.5 50.7 62.3 80.6 63.0 65.3 60.7 66.9 62.0 34.9 64.4
MaskCLIP++ 38.4 56.4 67.0 85.2 67.8 71.0 67.9 68.6 74.7 50.3 65.5
Table 1: Comparison of the classification accuracy of ground-truth masks (maskAcc) with different methods on multiple datasets. The same architecture (ViT-L/14) and input resolution (short side length 392) are used for all methods. “General”, “Earth”, “Medical”, “Engineer”, and “Agriculture” represent mean mask accuracy across five domain datasets in the MESS benchmark [4], with detailed scores for each individual dataset presented in the supplementary material.
Method CLIP arch. Training dataset A-847 PC-459 A-150 PC-59 PAS-20
Non-mask-based OVS method
OVSeg [35] ViT-L/14 COCO Stuff 9.0 12.4 29.6 55.7 94.5
CLIPSelf [65] ViT-L/14 COCO Stuff 12.4 - 34.5 62.3 -
SED [68] ConvNeXt-L COCO Stuff 13.7 22.1 35.2 60.6 96.1
CAT-Seg [11] ViT-L/14 COCO Stuff 16.0 23.8 37.9 63.3 97.0
Mask-based OVS method
MaskCLIP [15] ViT-L/14 COCO Panoptic 8.2 10.0 23.7 45.9 -
MasQCLIP [74] ViT-L/14 COCO Panoptic 10.7 18.2 30.4 57.8 -
ODISE [71] ViT-L/14 COCO Panoptic 11.1 14.5 29.9 57.3 -
SAN [73] ViT-L/14 COCO Stuff 12.4 15.7 32.1 57.7 94.6
SCAN [38] ViT-L/14 COCO Stuff 14.0 16.7 33.5 59.3 97.2
FC-CLIP [76] ConvNeXt-L COCO Panoptic 14.8 18.2 34.1 58.4 95.4
MAFT+ [27] ConvNeXt-L COCO Stuff 15.1 21.6 36.1 59.4 96.5
MaskCLIP++ w/ FC-CLIP ViT-L/14 COCO Stuff 15.4+0.6 21.3+3.1 37.1+3.0 62.6+4.2 96.4+1.0
MaskCLIP++ w/ MAFT+ ViT-L/14 COCO Stuff 16.8+1.7 23.9+2.3 38.2+2.1 62.5+3.1 96.8+0.3
Table 2: Comparisons with previous methods on open-vocabulary semantic segmentation task.
4.2 Main results
Mask classification. Tab. 1 presents the mask classification accuracy of CLIP models from different sources when processing input images and ground-truth masks at identical resolutions. In addition to the two pre-trained CLIP variants, we also demonstrate the performance of CLIP fine-tuned on COCO through distinct approaches: CLIPSelf [65] learned via patch-level feature distillation and CAT-Seg [11] trained with pixel-level cost aggregation. During inference, CLIPSelf performs mask pooling on the dense features of the final layer. Given that CAT-Seg operates as a pixel-level segmentation model, we leverage its CLIP parameters while maintaining consistency with the baseline methods by employing identical mask classification methodology. By directly utilizing higher-quality masks to fine-tune CLIP while adhering to consistent alignment principles, MaskCLIP++ achieves significant improvements over baselines across multi-domain datasets, outperforming other fine-tuning approaches.
Open-vocabulary semantic segmentation. Tab. 2 demonstrates the mIoU of MaskCLIP++ when employing mask generators from FC-CLIP [76] or MAFT+ [26]. Compared to MAFT+, our approach achieves mIoU improvements of +1.7, +2.3, +2.1, +3.2, and +0.3 points on the A-847, PC-459, A-150, PC-59, and PAS-20 datasets, respectively. These results indicate that our method more effectively harnesses the potential of existing mask generators through better mask classification capabilities.
Open-vocabulary panoptic and instance segmentation. Tab. 3 demonstrates that our method can be effectively integrated with existing instance-level mask generators. Furthermore, the enhanced classification capabilities of MaskCLIP++ also improve the performance across these segmentation paradigms.
Method PQ SQ RQ AP
MaskCLIP [15] 15.1 70.5 19.2 6.2
MasQCLIP [74] 23.3 - - -
ODISE [71] 22.6 - - 14.4
FC-CLIP [76] 26.8 71.5 32.2 16.8
MAFT+ [27] 27.1 73.5 32.9 -
MaskCLIP++ w/ FC-CLIP 27.7 72.0 33.6 17.3
MaskCLIP++ w/ MAFT+ 28.1 74.0 34.7 -
Table 3: Comparisons with previous methods on open-vocabulary panoptic and instance segmentation on A-150.
4.3 Ablation Studies
The ablation studies in Tab. 4 and Tab. 5 are conducted using ViT-B/16 models fine-tuned on COCO Panoptic [36], whereas those presented in Tab. 6 and Tab. 7 employed ViT-L/14 models fine-tuned on COCO Stuff [5].
Importance of consistency alignment. We use different parameterized similarity modeling methods to illustrate the importance of the consistent alignment principle. Tab. 4 elucidates the relationship between consistency alignment and overfitting through PSM application analysis during inference. We evaluate alignment consistency by toggling PSM usage after fine-tuning, while partitioning validation set categories into seen (S) and unseen (U) subsets to compute mIoU(S) and mIoU(U) respectively. Severe overfitting is identified when optimized mIoU(U) significantly underperforms its no-fine-tuning baseline. The table substantiates that the $\operatorname{Linear}\langle E_{m},E_{t}\rangle$ approach preserving consistency alignment achieves simultaneous improvements in both mIoU(S) and mIoU(U), demonstrating effective generalization maintenance.
PSM Use mIoU(S) mIoU(U) mIoU
No fine-tune - 22.6 12.8 17.0
$\langle\operatorname{Linear}(E_{m}),E_{t}\rangle$ 36.6 21.7 28.0
40.4 6.2 20.8
$\langle E_{m},\operatorname{Linear}(E_{t})\rangle$ 36.1 21.4 27.7
41.2 6.6 21.4
$\operatorname{Linear}\langle E_{m},E_{t}\rangle$ 44.8 25.1 33.5
45.3 25.3 33.8
Table 4: The phenomenon of (in)consistent alignment when fine-tuning with different PSMs. The PSMs prone to overfitting (red) paradoxically shows better performance when omitting the PSM during inference, revealing the occurrence of inconsistent alignment. In contrast, the blue PSM maintains consistent alignment while achieving comparable performance both with and without the PSM application, thus effectively mitigating overfitting risks.
Impact of prior quality during training. We refer to the regional information fed into Fuser of CLIP-V as “priors” and investigate how the quality of these priors affects fine-tuning performance. We experiment with pixels, generated masks that match GT masks according to IoU, GT masks, and the GT boxes as their minimum bounding rectangles. As shown in Tab. 5, using GT masks can achieve the best performance. Using the generated masks as prior yields lower performance compared to the GT masks, indicating that incorporating the mask generator during training actually hinders CLIP fine-tuning. Both GT masks and GT boxes priors outperform pixel priors, suggesting that fine-tuning CLIP from global recognition to region-level recognition is less challenging than fine-tuning it to pixel-level recognition.
Prior Type mIoU(S) mIoU(U) mIoU maskAcc
Pixels 43.2 21.4 30.6 53.3
Gen. masks 43.8 23.8 32.3 52.2
GT boxes 44.4 24.2 32.8 57.6
GT masks 45.3 25.3 33.8 58.3
Table 5: Impact of prior quality during training. “Gen. masks” denotes the matched generated masks. “GT boxes” denotes the minimum bounding rectangles of GT masks.
Impact of different mask generators. Except for FC-CLIP and MAFT+ in Tab. 2, various types of mask generators can be integrated with MaskCLIP++, as shown in Tab. 6. We use mask generators from Mask2Former [10] and FC-CLIP [76], which are typically considered as closed-vocabulary and open-vocabulary mask generators, respectively. These mask generators are all trained on COCO Panoptic [36] and evaluated on ADE20K [80] using mIoU. As shown in Tab. 6, the original performance of “open-vocabulary” mask generators do do not significantly outperform “closed-vocabulary” ones because of the “mis-alignment” as shown in Fig. 2(a). And using our MaskCLIP++ for mask classification can significantly improve the segmentation performance of all mask generators. This suggests that both open-vocabulary and closed-vocabulary mask generators possess an inherent category-independent mask generation capability that has yet to be fully harnessed.
Generator OV Gen. Only w/ MaskCLIP++
Mask2Former (T) 15.7 35.6
Mask2Former (B) 16.8 35.5
FC-CLIP (B) 19.1 36.8
FC-CLIP (L) 20.8 37.1
Table 6: Impact of different mask generators on OVS performance. “OV” indicates whether the mask generator is open-vocabulary. “Gen. Only” indicates the mIoU of the mask generator on A-150. “w/ MaskCLIP++” indicates the mIoU of the mask generator when integrated with MaskCLIP++.
Impact of training data size. Since the goal of our fine-tuning is to transfer CLIP’s open-vocabulary recognition capabilities to mask regions, rather than learning category-specific pixel groupings from segmentation datasets, our method can effectively fine-tune using a smaller amount of image segmentation data. We randomly sampled varying proportions of data from the COCO-Stuff training set for fine-tuning, using the same configuration as the model in Tab. 2. As shown in Tab. 7, even with only 1% of the data (approximately 1K images with annotations), the fine-tuned model still demonstrates competitive performance.
Data size A-847 PC-459 A-150 PC-59 Stuff
100% 38.4 56.4 67.0 85.2 67.8
10% 39.4 56.2 67.0 84.0 67.4
1% 38.7 55.4 66.3 83.5 65.5
0.1% 36.3 54.5 61.2 75.5 55.7
Table 7: Mask accuracy when fine-tuning with different amounts of data. “data size” indicates the proportion of data used for fine-tuning relative to the original COCO-Stuff training set.
Image                           CAT-Seg                        MAFT+                     MaskCLIP++ GT
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 5: Visualizations of open-vocabulary semantic segmentation on ADE20K [80]. Our method produces more complete masks than CAT-Seg and fewer biases in mask classification than MAFT+.
5 Discussion
Qualitative comparison. Fig. 5 presents qualitative comparisons of CAT-Seg [11], MAFT+ [27], and MaskCLIP++ on open-vocabulary semantic segmentation tasks. Compared to CAT-Seg, mask-based methods often produce higher-quality masks. Compared to MAFT+, our method retains more of the original CLIP’s knowledge without generating too many erroneous biases. Our method also predicts relatively correct semantics for some objects that are not annotated in the GT.
Segmentation without mask generators. Because the goal of our fine-tuning is to enhance the mask classification ability of CLIP, which indirectly enhances the local representation between different patch representations of the CLIP. Diverging from Sec. 3.2, we apply mask generator-free inference schemes on MaskCLIP++ to demonstrate the enhanced inter-patch relationships, as shown in Tab. 8.
Method Load PC-459 A-150 PC-59 Stuff Citys
SCLIP [59] 7.7 16.1 34.2 22.4 32.2
9.5 19.3 40.8 29.0 35.9
ClearCLIP [29] 7.9 16.7 36.0 23.9 30.0
9.6 19.2 40.8 30.4 33.5
ProxyCLIP  [30] 8.4 20.2 39.1 26.5 38.1
9.9 23.2 44.6 34.2 42.4
Table 8: Performance improvements achieved when combined with several unsupervised open-vocabulary semantic segmentation methods that do not rely on a mask generator. “Load” means whether to load model weights from MaskCLIP++. OpenAI CLIP ViT-B/16 are used here.
Additional computation beyond mask generators. Tab. 9 presents the computational costs of different modules in our model from Tab. 2 under varying numbers of proposal masks. The CLIP-V adopts the ViT-L/14 architecture with an input resolution of $392^{2}$, while the mask generator employs the ConvNeXt-L architecture operating at $800^{2}$ resolution. Using 150 object categories, the table reveals that the primary computational bottleneck remains in the mask generator since CLIP-V for mask classification typically requires lower resolutions. Furthermore, the Extractor and mask generator can execute computations in parallel during implementation, thereby reducing overall processing time.
Q CLIP-V PSM Mask Generator
Extractor Fuser Backbone Decoder
100 416.26 67.05 0.05 877.63 232.12
250 416.26 78.62 0.12 877.63 302.35
Table 9: FLOPs (G) of each module.
6 Conclusion
This paper addresses key challenges in open-vocabulary segmentation (OVS) by proposing MaskCLIP++, a high-quality mask tuning framework for CLIP that significantly simplifies the training paradigm. Our method achieves notable improvements in mask classification accuracy across multiple domain datasets, and when integrated with existing mask generators, demonstrates superior open-vocabulary image segmentation performances.
Supplementary Material for “High-Quality Mask Tuning Matters for Open-Vocabulary Segmentation”
Appendix A Method Details
A.1 Extracting mask embeddings of different architectures
The main text exclusively details mask embedding extraction from ViT-architecture CLIP models. This methodology can be extended to other CLIP architectures through the following adaptations: For architectures containing attention pooling layers without Transformer blocks, the Fuser module operates starting from the attention pooling layer, with weight functions identical to Eq. 2. In models employing global average pooling without Transformer components, the Fuser initiates computation from the global pooling layer, where the weight function degenerates to L1 normalization to implement masked average pooling.
A.2 Phenomenon of inconsistency alignment
We use a toy example to illustrate the potential consequences of destroying alignment consistency. Assume that we use the parameter $\Theta$ to update the mask embeddings, changing the original similarity from $s$ to $r$, and then optimize towards the target. However, due to the lack of explicit constraints on $\Theta$, there may be a situation as shown in Fig. 6, where the order of $s$ and $r$ is inconsistent, resulting in correct optimization on $r$ but incorrect optimization on $s$. For simplicity, $\mathcal{L}$ in Fig. 6 denotes an L1 loss without softmax, and only mask embeddings $m$ are updated in one round of gradient descent.
A.3 Mask classification score integration
The mask generator demonstrates inherent capability to produce classifications within its training vocabulary. When unambiguous semantic correspondence exists during inference, following established protocols in mask-based open-vocabulary segmentation (OVS) works [72, 71, 76], our method’s output class scores can be adaptively aggregated across the mask generator’s training vocabulary. This enables category probability estimation while maintaining compatibility with existing OVS frameworks.
Specifically, let $\mathcal{K}$ be the training vocabularies of the mask generator. $P_{c}$ and $P_{s}$ are the class probabilities of /Ours and the mask generator, respectively. The integrated class probabilities $P_{\gamma}$ can be expressed as follows:
$P_{\gamma}^{(i)}=\begin{cases}{P_{s}^{(i)}}^{\gamma}\cdot{P_{c}^{(i)}}^{(1-\gamma)},&\text{if}~{}i\in\mathcal{K}\\ P_{c}^{(i)}.&\text{otherwise}\\ \end{cases}$ (3)
We set $\gamma=0.1$ for our model when integrating with the mask generator. However, as demonstrated in Tab. 15, our’s performance enhancement stems primarily from improved mask classification capability rather than this integration strategy.
Appendix B Dataset Details
Dataset #Classes #Images
Train
COCO Stuff [5] 171 118280
COCO Panoptic [36] 133 118287
Eval
COCO Stuff 171 5000
COCO Panoptic 133 5000
ADE20K [80] 150 / 847 2000
PASCAL Context [43] 59 / 459 5000
PASCAL VOC [17] 20 (w/o background) 1449
Cityscapes [13] 19 500
Eval (MESS - General)
BDD100K [75] 19 1000
Dark Zurich [49] 20 50
MHP v1 [32] 19 980
FoodSeg103 [66] 104 2135
ATLANTIS [16] 56 1295
DRAM [12] 12 718
Eval (MESS - Earth)
iSAID [64] 16 4055
ISPRS Potsdam [56] 6 504
WorldFloods [24] 3 160
FloodNet [48] 10 5571
UAVid [41] 8 840
Eval (MESS - Medical)
Kvasir-Inst. [45] 2 118
CHASE DB1 [18] 2 20
CryoNuSeg [42] 2 30
PAXRay-4 [50] 2 180
Eval (MESS - Engineer)
Corrosion CS [3] 4 44
DeepCrack [37] 2 237
ZeroWaste-f [2] 5 929
PST900 [53] 5 288
Eval (MESS - Agriculture)
SUIM [23] 8 110
CUB-200 [58] 201 5794
CWFID [21] 3 21
Table 10: Statistics of datasets used in our experiments.
Tab. 10 provides detailed statistics for the datasets used in our experiments. Tab. 11 lists the categories in ADE20K that are present in COCO Panoptic (seen) or absent from COCO Panoptic (unseen). This split is used in our experiments to evaluate the generalization capability in Sec. 4.3.
Split Name
Seen (64) wall, building, sky, floor, tree, ceiling, road, bed, windowpane, grass, cabinet, sidewalk,
person, door, table, mountain, plant, curtain, chair, car, water, sofa, shelf, sea,
mirror, rug, fence, rock, lamp, counter, sand, sink, refrigerator, stairs, pillow, river,
bridge, toilet, flower, book, bench, palm tree, boat, bus, towel, light bulb, truck,
television receiver, airplane, apparel, bottle, tent, oven, food, microwave, plant pots,
animal, bicycle, blanket, vase, traffic light, plate, cup, clock
Unseen (86) earth, painting, house exterior, field, armchair, seat, desk, wardrobe,
bathtub, railing, cushion, pedestal, box, column, signboard, chest of drawers, skyscraper,
fireplace, grandstand, path, runway, case, pool table, screen door, stairway, bookcase,
window screen, coffee table, hill, countertop, stove, kitchen island, computer, swivel chair,
arcade machine, hovel, tower, chandelier, awning, streetlight, booth, dirt track, pole, land,
bannister, escalator, ottoman, buffet, poster, stage, van, ship, fountain, conveyer belt,
washer, plaything, swimming pool, stool, barrel, basket, waterfall, bag, minibike, cradle,
step, tank, trade name, lake, dishwasher, projection screen, sculpture, exhaust hood, sconce,
tray, ashcan, ceiling fan, pier, crt screen, monitor, bulletin board, shower, radiator,
canopy, flag, bar, ball
Table 11: Categories in ADE20K that are present in COCO Panoptic (seen) or are absent from COCO Panoptic (unseen).
Method BDD100K Dark Zurich MHPv1 FoodSeg103 ATLANTIS DRAM
OpenAI CLIP 44.6 27.5 48.6 57.1 58.1 89.3
EVA02 CLIP 44.7 34.3 49.2 61.4 61.0 91.0
CLIPSelf 51.8 30.6 61.0 47.4 59.1 65.3
CAT-Seg 59.3 40.3 49.8 54.6 65.3 94.6
MaskCLIP++ 67.2 48.3 64.9 64.7 64.5 98.0
(a) General
Method iSAID ISPRS Potsdam WorldFloods FloodNet UAVid
OpenAI CLIP 59.2 65.5 65.9 65.9 48.9
EVA02 CLIP 61.0 68.4 71.4 59.6 42.2
CLIPSelf 43.6 41.7 42.4 38.3 62.6
CAT-Seg 59.8 76.9 71.0 62.3 64.4
MaskCLIP++ 60.5 80.0 75.1 63.0 64.4
(b) Earth Monitoring
SUIM CUB-200 CWFID
55.5 54.7 42.8
57.6 60.0 39.7
56.0 51.1 61.9
72.9 47.2 73.0
73.2 40.8 82.5
(c) Agriculture and Biology
Method Kvasir Inst. CHASE DB1 CryoNuSeg PAXRay-4
OpenAI CLIP 93.2 50.0 50.0 50.8
EVA02 CLIP 95.0 50.0 50.0 51.7
CLIPSelf 72.0 50.0 50.0 50.0
CAT-Seg 89.8 50.0 50.0 58.1
MaskCLIP++ 99.2 50.0 50.0 99.4
(d) Medical Sciences
Corrosion CS DeepCrack ZeroWaste-f PST900
25.5 69.2 32.5 19.4
25.0 58.9 33.0 18.4
31.6 76.8 45.2 27.5
21.6 61.0 30.2 26.7
24.4 85.7 57.4 33.8
(e) Engineering
Table 12: Mask accuracy on MESS datasets [4].
Appendix C Additional Experiments
C.1 Mask accuracy on MESS
Tab. 12 complements Tab. 1 by listing the mask accuracy of different models on each dataset in MESS [4]. MaskCLIP++ achieves better maskAcc on 17 out of 22 datasets, providing compelling empirical evidence that the proposed method significantly enhances mask classification accuracy while simultaneously preserving the inherent generalization capabilities of the original CLIP.
However, note that in specialized fine-grained classification tasks like the bird-oriented CUB-200 dataset [58], where category labels are semantically close and images usually have a single foreground subject, fine-grained category domain-specific knowledge impacts performance more than localized perception. Thus, CLIP models fine-tuned on COCO generally perform worse than the original CLIP in such cases.
C.2 Effectiveness of mask tuning on different CLIP architectures
To validate the generalizability of our approach across varying CLIP architectures, we conduct fine-tuning experiments on the COCO Panoptic dataset [36] using distinct CLIP architectures and report the mask classification accuracy of our method as documented in Tab. 13.
Arch. A-847 PC-459 A-150 PC-59 Stuff
RN50x16 24.1 37.5 50.4 73.3 54.1
ConvNeXt-B 30.0 42.1 56.4 76.3 55.6
ConvNeXt-L 33.0 50.8 61.0 78.7 59.0
ConvNeXt-XXL 38.7 56.1 64.6 81.5 62.5
ViT-B/16 29.3 45.7 58.3 78.6 58.0
ViT-L/14 37.5 56.9 66.0 83.2 63.6
Table 13: Mask accuracy of MaskCLIP++ on different architectures fine-tuned on COCO Panoptic [36].
C.3 Other ablation studies
Effect of Fuser depth. The positioning of mask insertions exerts non-negligible effects on model performance. As delineated in Sec. 3.2, the quantity of Transformer blocks in the Fuser module can be mathematically formulated as $L-K$. Tab. 14(a) presents a systematic analysis of how varying Fuser depths impact OVS performance. Considering the computational overhead that scales linearly with mask quantity when increasing Fuser depth, we empirically select $L-K=2$ as the default configuration.
$\bm{L-K}$ mIoU PQ
1 33.3 24.1
2 33.5 24.5
3 33.8 24.4
4 33.5 24.4
(a) Effect of Fuser depth.
$\bm{\alpha}$ mIoU
$e^{5}$ 33.1
$e^{-5}$ 33.5
learnable 33.8
(b) Effect of $\alpha$.
Table 14: Ablation studies on the effect of Fuser depth and $\alpha$.
Effect of $\alpha$. As defined in Eq. 2, the bigger $\alpha$ is, the less the semantic similarity of the image context is utilized. As $\alpha\rightarrow+\infty$, $E_{m}^{(l+1)}$ degenerates to the mask average pooling result of $F^{(l)}$, thereby inadequately leveraging contextual information beyond the masked regions. Conversely, when $\alpha=0$, $E_{m}^{(l+1)}$ reduces to the CLS token after $l$-th layer. As shown in Tab. 14(b), selecting a smaller $\alpha$ value yields superior performance. Therefore, we initialize $\alpha$ as $e^{-5}$ and adaptively adjust it in the logarithmic space during training.
Effect of classification ensemble. To demonstrate that the performance gains of our method primarily stem from enhanced mask classification capabilities, we present in Tab. 15 the comparative analysis of implementing versus omitting the ensemble strategy.
Method Ensemble A-847 PC-459 A-150 PC-59
FC-CLIP 12.1 12.7 27.4 42.8
14.8 18.2 34.1 58.4
MaskCLIP++ 15.3 21.4 36.8 62.4
15.4 21.3 37.1 62.6
Table 15: Effect of classification ensemble.
Image                          Semantic Segmentation                Panoptic Segmentation Instance Segmentation
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 7: More visualizations of semantic- and instance-level of open-vocabulary segmentation. We use mask generator from FC-CLIP [76] with ConvNeXt-L backbone. And use CLIP ViT-L/14 fine-tuned on COCO Panoptic [36]. Images are from ADE20K [80]. The confidience threshold for instance segmentation is 0.3.